OpenStudy (anonymous):

Trevor solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x - 2y = 10 y = 5 - 2x x - 2(5 - 2x) = 10 x - 10 + 4x = 10 5x - 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 Answer He made no mistake He subtracted 10 from the right side instead of adding 10 to the right side He combined like terms incorrectly, it should have been 4x instead of 5x He should have substituted 5 + 2x

OpenStudy (anonymous):

He subtracted 10 from the right side instead of adding 10 to the right side

OpenStudy (lalaly):

5x - 10 = 10 5x = 0 x = 0 it should be 5x=10+10 5x=20 x=4